### v3

# Soliton resuscitations: asymmetric revivals of the breathing mode of an atomic bright soliton in a harmonic trap

Abstract
An atomic bright soliton realised in a quasi-one-dimensional Bose-Einstein condensate can be considered as an open quantum system. The soliton’s breathing mode, for example, is damped by emission of atoms from the soliton to spatial infinity, which thus acts as a Markovian environment for the soliton. If the soliton is held in a shallow harmonic trap, however, the environment becomes non-Markovian: emitted atoms oscillate in the trap and eventually return to the soliton, interfering with it, producing periodic revivals of the breathing mode (“resuscitations”). The amplitude envelopes of these breathing revivals shows a curious asymmetry, with a gradual increase in breathing amplitude followed by sudden drop in amplitude that becomes more and more pronounced in later revivals. We explain this asymmetrical revival pattern in the non-Markovian revivals by deriving a close analytical approximation to the Bogoliubov-de Gennes frequency spectrum for the weakly trapped soliton.
I I. Introduction
I.1 I.A. Bright solitons
Bright solitons are named for their appearance as self-focused light pulses in nonlinear optical fibers[1], but they can also occur in condensed Bose gases[2, 3], as self-binding concentrations of atomic vapor rather than of light. Similar solitons occur in both kinds of medium, because bright solitons are solutions to the non-linear Schrödinger equation in one spatial dimension with a negative nonlinear term. When the nonlinear Schrödinger equation appears as the Gross-Pitaevskii equation for the macroscopic wave function of a dilute Bose-Einstein condensate, a negative scattering length for interatomic collisions provides the necessary negative nonlinearity for bright solitons. In order to support such solitons, the macroscopic wave function $\psi(x,t)$ of the condensate must effectively be confined to only one spatial dimension.
Experiments have produced such atomic bright solitons in Lithium[2, 3, 4], Potassium[5], and Cesium[6] vapors. Further experiments have observed reflections[7] and collisions[8] of atomic bright solitons. An atomic Mach-Zehnder interferometer using bright solitons has been shown to have enhanced fringe visibility[9]. Trains of multiple bright solitons[3], and collective motional modes of the trains[10], have also been studied.
I.2 I.B. Collective “breathing” excitations
Collective excitations of individual solitons have attracted interest as well, and indeed the distinction between collective motional modes of multi-soliton groups, and the vibration-like excitation modes of individual solitons, is not really sharp. Higher-order “breather” solitons[11, 12, 13, 14] can be considered single solitons with nonlineary large-amplitude breathing excitations, or they can be seen as interactions of two or more solitons that repeatedly merge and separate. Quantum effects in these nonlinear excitations, for the dilute condensed Bose gas beyond the mean-field nonlinear Schrödinger equation, have been studied theoretically using truncated Wigner[11], Bethe Ansatz[12] and Bogoliubov-de Gennes theory[13].
The low-amplitude, linear limit of a single-soliton breathing mode, as illustrated in Fig. 1, has also been induced in experiments[15], with the breathing clearly observed over many periods.
Figure 1: Small-amplitude soliton breathing in mean field theory, with no trapping potential. The gas density in the attractively interacting condensate is self-bound in a $\mathrm{sech}^{2}$ profile; the width and peak density of the spatial profile oscillate periodically in time, while the oscillation amplitude decays due to dispersion.
Figure 2: Asymmetrical resuscitations of soliton breathing in a harmonic trap of frequency $\omega$, as shown in the peak condensate density $|\psi(0,t)|^{2}$ as a function of time in trap periods. The evolution is given by the one-dimensional Gross-Pitaevskii equation from the non-stationary initial wave function $\psi(x,0)=A(\tilde{\kappa}/\sqrt{g})\mathrm{sech}(\tilde{\kappa}x)$ for $\tilde{\kappa}=\sqrt{50}a_{0}^{-1}$, where $a_{0}=\sqrt{\hbar/(2M\omega)}$ (with atomic mass $M$) is the harmonic trap width. The case $A=1$ would be a perfect bright soliton, with no breathing; in this case $A=1.01$, providing a small-amplitude breathing excitation at the chemical potential frequency $\hbar\kappa^{2}/(2M)=A^{4}(\tilde{\kappa}a_{0})^{2}\omega\doteq 52\omega$, while the breathing amplitude revives every half-period of the harmonic trap. The first revival begins noticeably sooner than a trap half-period. Note the “trumpet-shaped” asymmetrical envelopes of the revived oscillations.
This linear breathing is not a unique normal mode, however. Small-amplitude excitations of a single bright soliton with no external trapping potential have a continuous frequency spectrum[16], and half of all modes (the ones with even spatial parity) include some component of “breathing” (oscillation in the soliton width and amplitude). Although there is thus no unique breathing mode, and no unique breathing mode frequency, typical initial state preparation mainly excites low-frequency perturbations of the background soliton, and interference between the background order parameter and low-frequency perturbations makes the gas density breathe at the chemical potential frequency, while dispersion in the excitation continuum makes the breathing amplitude gradually decay, within a $t^{-1/2}$ envelope[17], as is visible in Fig. 1.
For a breathing soliton which is purely self-trapped, with no external potential, part of this dispersive relaxation process includes emission of small-amplitude waves to infinity. For atomic bright solitons described in Gross-Pitaevskii mean field theory, this implies ejection of atoms from the soliton, which carry away energy as well as particles from it forever. Infinity—or simply spatial regions well outside the soliton’s width—thereby act as an environment within which the soliton moves or breathes as an open system.
I.3 I.C. Breathing solitons in traps
Bright solitons in experiments, however, are generally held in trapping potentials that can be approximated well as harmonic. Ejected atoms will therefore not escape to infinity, but rather oscillate in the trap and return to the soliton. If they retain phase coherence with the macroscopic wave function of the soliton, these returning atoms will interfere with the soliton, and the breathing mode will therefore revive periodically. Fig. 2 shows an example of such breathing revivals—“resuscitations”—as predicted by mean field theory (i.e. under the time-dependent Gross-Pitavskii nonlinear Schrödinger equation in one spatial dimension).
These resuscitations of the soliton breathing mode are thus a simple example of non-Markovian open-system evolution, in which the environment effectively remembers information that has been received from the system, so that its present effects on the system depend on the system’s prior history. In this case the non-Markovian nature of the environment is simply oscillation of atoms in a harmonic potential, and so the most striking feature of the resuscitations is simply that they recur at every half-period of the harmonic trap. Looking more closely at Fig. 2, however, we notice the curiously asymmetric, trumpet-shaped envelopes of these resuscitations, which seem to become more and more asymmetric in each subsequent revival. Even such a simply non-Markovian environment as a harmonic trap can produce complex open-system evolution, through interplay between the non-Markovian environment and non-trivial system dynamics, such as the breathing of a bright soliton with attractive mean-field interactions.
In this paper we analytically study these temporally asymmetrical soliton breathing revivals, by solving the linearized mean-field evolution of a soliton in a shallow harmonic trap using matched asymptotics in the Bogoliubov-de Gennes equations. Our results illustrate the complexity of non-Markovian evolution even in a simple instance, and they will also provide a mean-field background against which quantum many-body effects may be discerned in future studies and experiments.
I.4 I.D. Paper structure
Our paper is structured to avoid burdening readers with technical details before the meaning of the details is clear. In Section II we briefly review the untrapped atomic bright soliton in Gross-Pitaevskii mean field theory, as well as its linearized Bogoliubov-de Gennes perturbations. In Section III we then briefly present our results for the modified Bogoliubov-de Gennes modes in the harmonic trap, without derivation, and in Section IV we show how these results lead to the asymmetrical resuscitations that we have described, before concluding and discussing the outlook for future study in Section V. Technical details, including the derivation by matched asymptotics of the main results that we applied in Section III, are then provided in four rather lengthy appendices.
II II. The Untrapped Bright Soliton
II.1 II.A. The mean field background
We consider a quasi-one-dimensional Bose-Einstein condensate with attractive interactions, using Gross-Pitaevskii mean field theory with Bogoliubov-de Gennes linearized excitations. As we note in Appendix A, mean field theory can be expected to be highly accurate for slightly excited atomic bright solitons, even with modest condensate numbers, because the quantum depletion
$\langle\hat{\psi}^{\dagger}(x)\hat{\psi}(x)\rangle-|\Psi_{0}|^{2}$ (1)
of a stationary atomic bright soliton is only one-third: not one-third of the condensate, but one-third of an atom. As the zeroth-order background state of the condensate’s macroscopic wave function $\psi(x,t)$, we therefore assume the bright soliton form
$\Psi_{0}(x,t)=e^{i\frac{\hbar\kappa^{2}}{2M}t}\frac{\kappa}{\sqrt{g}}\mathrm{sech}(\kappa x)\;,$ (2)
where g-g is the one-dimensional scattering strength for collisions between the condensate particles, $M$ is the particle mass, and $\kappa$ is a constant inverse length. In one dimension, the scattering strength $g$ is an inverse length; for quasi-one-dimensionally trapped gases, $g$ is proportional to the three-dimensional scattering length divided by the mean cross section of transverse confinement. Assuming here negative scattering length, we consider only $g>0$.
The background $\Psi_{0}$ of (2) solves the Gross-Pitaevskii nonlinear Schrödinger equation, which with our definition of $g$ reads as
$i\hbar\frac{\partial}{\partial t}\psi=-\frac{\hbar^{2}}{2M}\frac{\partial^{2}}{\partial x^{2}}\psi+V(x)\psi-\frac{\hbar^{2}g}{M}|\psi|^{2}\psi\;$ (3)
in general; the solution $\psi\to\Psi_{0}$ as given by (2) applies exactly in the special case $V(x)=0$, but also offers a good approximate solution for any $V(x)$ that varies slowly over the soliton width scale $\kappa^{-1}$.
The inverse length $\kappa$ is arbitrary, inasmuch as it depends on how many atoms are in the Bose-condensed gas. Conversely, for given $\kappa$ the condensate atom number in the bright soliton is
$N_{0}=\int\!dx\,|\Psi_{0}|^{2}=\frac{\kappa}{g}\int\!d(\kappa x)\,\mathrm{sech}^{2}(\kappa x)=\frac{2\kappa}{g}\;.$ (4)
II.2 II.B. Bogoliubov-de Gennes perturbations
For small, time-dependent perturbations around the bright soliton background,
$\psi(x,t)=\Psi_{0}(x,t)+e^{i\frac{\hbar\kappa^{2}}{2M}t}\delta\psi(x,t)\;,$ (5)
the perturbation $\delta\psi$ can be expanded in a set of normal modes:
$\delta{\psi}(x,t)=$ $\int_{-\infty}^{\infty}\!dk\,\Big{[}{\beta}_{k}e^{-i\Omega(k)t}u_{k}(x)+{\beta}^{*}_{k}e^{i\Omega(k)t}v^{*}_{k}(x)\Big{]}$
$\Omega(k)=$ $\frac{\hbar}{2M}(k^{2}+\kappa^{2})$ (6)
for arbitrary (small) complex constants $\beta_{k}$. In addition to this continuous set of $(u_{k},v_{k})$ excitations, the bright soliton also possesses two discrete zero modes; as we discuss in Appendix B, these modes are not directly involved in the breathing excitations that we discuss in this paper. One of them is indirectly involved, in that the requirement to “fix” this zero mode will resolve an ambiguity in distinguishing which part of our initial $\psi(x,0)$ is the soliton background and which part is the perturbation.
To make this perturbed $\psi(x,t)$ solve the Gross-Pitaevskii equation (3) to first order in $\delta\psi$, particular functional forms are required for the mode functions $u_{k}(x)$ and $v_{k}(x)$. These forms are known explicitly and exactly for the case $V=0$ [16], and it will be a major goal in this paper to find them approximately for a harmonic $V$. For $V=0$ we have the normalized mode functions are[16]
$u_{k}(x)$ $=\frac{1}{\sqrt{2\pi}}\frac{\kappa^{2}}{k^{2}+\kappa^{2}}\left(\frac{k^{2}-\kappa^{2}}{\kappa^{2}}+\frac{2ik}{\kappa}\tanh(\kappa x)+\mathrm{sech}^{2}(\kappa x)\right)e^{ikx}$
$v_{k}(x)$ $=\frac{1}{\sqrt{2\pi}}\frac{\kappa^{2}}{k^{2}+\kappa^{2}}\mathrm{sech}^{2}(\kappa x)e^{ikx}\;.$ (7)
In this paper we will analyse only the particular kind of small perturbations around $\Psi_{0}$ that evolve from slightly stretched or compressed initial conditions of the form
$\psi(x,0)=(1+\delta)\frac{\tilde{\kappa}}{\sqrt{g}}\mathrm{sech}(\tilde{\kappa}x)=\Psi_{0}(x,0)+\delta\psi(x,0)\;,$ (8)
where $\delta$ is small, and where there is a subtle distinction between the $\tilde{\kappa}$ scale that appears in our initial condition, and the corresponding scale $\kappa$ in our soliton background.
If $\delta>0$ then the initial wave function is slightly too high, in comparison to its width, to be a perfect $\Psi_{0}$ soliton with $\kappa\to\tilde{\kappa}$; if $\delta<0$ then $\psi(0,0)$ is not quite high enough. Our $\psi(x,0)$ is thus certainly not a perfect bright soliton, and we will consider $\psi(x,0)=\Psi_{0}(x,0)+\delta\psi(x,0)$. This leaves an ambiguity, however, in identifying which value of $\kappa$ we should use in $\Psi_{0}(x,0)$, since choosing a slightly different value of $\kappa$ would make only a small change in $\Psi_{0}$, and this small difference could simply be considered part of $\delta\psi(x,0)$.
As we show in Appendix B, however, requiring that $\delta\psi(x,t)$ should remain small over long times, and not grow slowly but steadily larger with $t$, selects one particular value for $\kappa$ for a given $\tilde{\kappa}$ and $\delta$:
$\kappa=(1+\delta)^{2}\tilde{\kappa}\;$. (9)
This provides the unambiguous initial perturbation
$\delta\psi(x,0)$ $=\frac{1}{1+\delta}\frac{\kappa}{\sqrt{g}}\mathrm{sech}\Big{(}\frac{\kappa x}{(1+\delta)^{2}}\Big{)}-\frac{\kappa}{\sqrt{g}}\mathrm{sech}(\kappa x)$
$=\delta\frac{\kappa}{\sqrt{g}}[2\kappa x\tanh(\kappa x)-1]\mathrm{sech}(\kappa x)+\mathcal{O}(\delta^{2})\;.$ (10)
II.3 II.C Breathing evolution
When our bright soliton is initially perturbed by a small $\delta\psi(x,0)$ of the form (II.2), then, we can use the orthonormality of the $u_{k}(x),v_{k}(x)$ of (II.2)
$\int_{-\infty}^{\infty}\!dx\,(u_{k}u_{k^{\prime}}^{*}-v_{k}v_{k^{\prime}}^{*})$ $=\delta(k-k^{\prime})$
$\int_{-\infty}^{\infty}\!dx\,(u_{k}v_{k^{\prime}}-v_{k}u_{k^{\prime}})$ $=0$ (11)
to obtain
$\beta_{k}$ $=\int_{-\infty}^{\infty}\!dx\,\Big{(}u_{k}^{*}(x)\delta\psi(x,0)-v_{k}^{*}(x)\delta\psi^{*}(x,0)\Big{)}$
$=-\sqrt{\frac{\pi}{2g}}\,\delta\,\mathrm{sech}\Big{(}\frac{\pi k}{2\kappa}\Big{)}\;,$ (12)
where the exact evaluation of the integral is obtained through integration by parts followed by contour integration. Note that $\beta_{-k}=\beta_{k}$, so that the reflection symmetry of our initial state is preserved through time evolution.
With this specific form (II.3) of $\beta_{k}$, we can integrate (VII.1) to obtain the breathing perturbation $\delta\psi(x,t)$ for all $t$. For small $\delta$, the density expanded to first order in $\delta$
$|\psi(x,t)|^{2}\doteq|\Psi_{0}|^{2}+\Psi_{0}\delta\psi^{*}+\Psi_{0}^{*}\delta\psi$ (13)
is an excellent approximation to the exact $|\psi(x,t)|^{2}$ over many breathing periods. A simple measure of the overall amplitude of the breathing mode is then $|\psi(0,t)|^{2}$, which from (VII.1), (II.2), and (II.3) is given by
$\frac{g}{\kappa^{2}}|\psi(0,t)|$ $=1-\delta\ B_{0}(t)+\mathcal{O}(\delta^{2})$
$B_{0}(t)$ $=\int_{-\infty}^{\infty}\!\frac{dk}{\kappa}\,\mathrm{sech}\Big{(}\frac{\pi k}{2\kappa}\Big{)}\cos\Big{(}\frac{\hbar(k^{2}+\kappa^{2})t}{2M}\Big{)}\;.$ (14)
For later times, the integral that defines the dimensionless breathing amplitude $B_{0}(\tau)$ in (II.3) can be evaluated using the method of steepest descents to reveal the asymptotic behavior
$B_{0}(t)=\sqrt{\pi}\frac{\cos\Big{(}\frac{\hbar\kappa^{2}}{2M}t+\frac{\pi}{4}\Big{)}}{\sqrt{\frac{\hbar\kappa^{2}}{2M}t}}+\mathcal{O}(t^{-3/2})\;.$ (15)
The breathing at the chemical potential frequency $\hbar\kappa^{2}/(2M)$ which we have described thus decays in amplitude as $t^{-1/2}$, due to dispersion.
Figure 1 shows an example of this breathing in the absence of a trapping potential, for an initial state (52) with $\delta=0.2$. The breathing that occurs from this initial state, under Gross-Pitaevskii mean-field evolution, does indeed have the frequency
$\omega_{B}=\frac{\hbar\kappa^{2}}{2M}=(1+\delta)^{4}\frac{\hbar\tilde{\kappa}^{2}}{2M}\;.$ (16)
In this illustrative case $\delta=0.2$ is small enough that the Bogoliubov-de Gennes linearization of the perturbation is still quite accurate, but the quartic $(1+\delta)^{4}\doteq 2.1$ factor in $\omega_{B}$ is large enough to confirm that the soliton itself really “knows” that $\kappa$, rather than $\tilde{\kappa}$, is its correct background soliton width: There are nearly twenty-one breathing periods in the time plotted, not just ten.
Because the Bogoliubov-de Gennes breathing oscillation dies away from dispersion, at very late times the post-Bogoliubov-de Gennes $\mathcal{O}(\delta^{2})$ corrections to $\psi$ may be more significant locally than the leading-order term (13). We can actually identify the main post-Bogoliubov-de Gennes effect, however, using only the leading order $\delta\psi(x,t)$ from (VII.1) with the coefficients $\beta_{k}$ from (II.3).
II.4 II.D Particle emission
The reason we can find $\mathcal{O}(\delta^{2})$ effects using only our $\delta\psi$ to order $\delta$ is that for $|x|\gg 1/\kappa$, the soliton background $\Psi_{0}(x,t)$ is exponentially small. In the region of large $x$, therefore, the full density is $|\delta\psi(x,t)|^{2}+\mathcal{O}(\delta^{3})$. We can thus compute the number of particles which have moved far away from the soliton at time $t$ as
$\Delta N(t)=\int_{L}^{\infty}\!dx\,\Big{(}|\delta\psi(x,t)|^{2}+|\delta\psi(-x,t)|^{2}\Big{)}$ (17)
for any $L\gg 1/\kappa$.
Evaluating $\delta\psi(x,t)$ for $|x|\geq L$ and at late $t$ using steepest descents, this lets us obtain
$\lim_{t\to\infty}\Delta N(t)=\delta^{2}N_{0}+\mathcal{O}(\delta^{3},t^{-1})\;.$ (18)
This lets us know, without actually calculating the $\mathcal{O}(\delta^{2})$ correction to $\delta\psi$, that the number of particles remaining in the soliton itself must gradually decrease over time, as the breathing mode decays. Mathematically we can recognize that the breathing mode decays due to dispersion, but the excitation energy is physically carried away from the soliton by particles that are emitted from the soliton to infinity.
This shaking loose of particles as the soliton breathes suggests the main difference which we can expect in the breathing of a soliton in a weak harmonic trap, compared to a soliton with no trap. As long as the trap is sufficiently weak, the soliton’s breathing oscillation will initially be the same as without the trap, and so will the initial stages of its $1/\sqrt{t}$ dispersive decay. In the trap, however, particles which are shaken loose from the oscillating soliton will not be able to escape to infinity: instead they will return to the soliton, after half a trap period, and interfere again with $\Psi_{0}$. The breathing mode will resuscitate.
III III. Collective modes of a bright soliton in a shallow trap
III.1 III.A: The shallow trap
Our goal is to understand how the breathing mode evolves when the condensate is held in a harmonic confining potential $V(x)\to\frac{M\omega^{2}}{2}x^{2}$. To maintain the open-system picture in which the soliton as a self-interacting system is only weakly or slowly affected by a long-range environment, we will focus on the limit in which the harmonic trap, with trap width $a_{0}=\sqrt{\hbar/(2M\omega)}$, is very broad and shallow compared to the bright soliton’s width $1/\kappa$:
$\frac{1}{\kappa a_{0}}=\frac{1}{\kappa}\sqrt{\frac{2M\omega}{\hbar}}=:\varepsilon\ll 1\;.$ (19)
In this limit the $V\to 0$ bright soliton $\Psi_{0}(x,t)$ is still a solution to the Gross-Pitaevskii equation with the harmonic trap $V\not=0$, up to corrections that are only of order $\varepsilon^{4}$, which we will consistently neglect throughout this paper.
III.2 III.B: Discrete spectrum
In this small-$\varepsilon$ limit the $u_{k}(x)$ and $v_{k}(x)$ modes of (II.2) are also still solutions to the Bogoliubov-de Gennes solutions, locally for $|x|\ll a_{0}$. Out of the continuum of $k$ modes that appear in the untrapped problem, however, only a particular discrete set of $k_{n}$ (for whole-number $n$) can be extended into solutions for all $x$. This will imply the discrete set of eigenfrequencies
$\delta\psi(x,t)$ $=e^{i\frac{\hbar\kappa^{2}}{2M}t}\sum_{n}\left(\beta_{n}e^{-i\Omega_{n}t}u_{n}(x)+\beta^{*}_{n}e^{i\Omega_{n}t}v_{n}^{*}(x)\right)$
$\Omega_{n}$ $=\frac{\hbar k_{n}^{2}}{2M}\;.$ (20)
The discrete set of modes will have spatial parity given by $n$,
$\left(\begin{matrix}u_{n}(x)\\ v_{n}(x)\end{matrix}\right)\underset{|x|\ll a_{0}}{\longrightarrow}Z_{n}\left[\left(\begin{matrix}u_{k_{n}}(x)\\ v_{k_{n}}(x)\end{matrix}\right)+(-1)^{n}\left(\begin{matrix}u_{-k_{n}}(x)\\ v_{-k_{n}}(x)\end{matrix}\right)\right]$ (21)
and will have finite normalization
$\int\!dx\,(u^{*}_{m}u_{n}-v^{*}_{m}v_{n})=\delta_{mn}$ (22)
ensured by the normalization factors $Z_{n}$.
III.3 III.C: Results from matched asymptotics
III.3.1 Discrete frequencies
In Appendix C we exploit the smallness of $1/\kappa a_{0}=\varepsilon$ to apply the method of matched asymptotics, and construct the complete set of discrete $(u_{n},v_{n})$ by smoothly matching $(u_{k},v_{k})$ solutions from (II.2) for $|x|\lesssim\varepsilon^{1/2}a_{0}$ onto parabolic cylinder functions for $|x|\gtrsim\varepsilon^{1/2}a_{0}=\varepsilon^{-1/2}\kappa^{-1}$. This lengthy analysis in Appendix C supplies the key results of our paper:
$k_{n}$ $=\varepsilon\kappa\sqrt{n+\frac{1}{2}+\Delta_{n}}$ (23)
$\Omega_{n}$ $=\frac{\hbar\kappa^{2}}{2M}+\omega\Big{(}n+\frac{1}{2}+\Delta_{n}\Big{)}$ (24)
$\Delta_{n}$ $=\frac{2}{\pi}\cos^{-1}\left(\frac{1-\varepsilon^{2}n}{1+\varepsilon^{2}n}\right)\;.$ (25)
Equation (25 says that $\Delta_{n}$ rises smoothly from 0 for small $n$ through 1 around $n\simeq\varepsilon^{-2}$, and ultimately approaches 2 for $n\to\infty$. The fact that $\Delta_{n\to\infty}=2$ means that $\nu_{N-2}=N$ for $N\to\infty$, so that the same large frequency range which holds $N\gg 1$ harmonic oscillator eigenfrequencies only holds N2N-2 of the bright soliton Bogoliubov-de Gennes eigenfrequencies.
The missing two eigenfrequencies are the zero mode and the Kohn dipole mode, which are discussed in Appendix B and whose frequencies lie below the $\hbar\kappa^{2}/(2M)$ gap. We could consider them to be the cases n=2n=-2 and n=1n=-1, respectively, since the zero mode is spatially even and the dipole mode is odd. Since the bright soliton amplitude $\kappa$ could in principle be raised continuously from zero, without ever altering the extremely high-frequency end of the excitation spectrum except by the overall chemical potential shift, the total number of eigenfrequencies below some high upper bound $\Omega_{N}=\hbar\kappa^{2}/(2M)+\omega(N+1/2)$, including the two modes below the $\hbar\kappa^{2}/(2M)$ gap, must be $N$ for all solitons. Pushing the zero mode and Kohn mode down below the gap means that the N2N-2 modes that remain above the gap are smoothly stretched out in frequency to cover the same total frequency range with two fewer eigenfrequencies.
This subtle distortion of the Bogolibov-de Gennes eigenspectrum, in comparison to the evenly spaced harmonic oscillator spectrum, will be responsible for the main effect that we discuss in this paper. The fact that all the even-mode eigenfrequencies $\Omega_{2n}$ are spaced nearly $2\omega$ apart is responsible for the revivals of soliton breathing after every half-period $\pi/\omega$. The fact that all the intervals between successive $\Omega_{2n}$ are actually not quite exactly $2\omega$, but instead are slightly and systematically stretched according to (25) and (24), is responsible for the increasingly assymetrical forms of the resuscitations that we saw in Fig. 2.
III.3.2 Normalization
In Appendix D we use our matched asymptotics solutions from Appendix C to determine the normalization factors $Z_{n}$, finding
$Z_{n}^{2}$ $=\frac{dk_{n}}{dn}\;,$ (26)
where we differentiate $k_{n}$ with respect to $n$ as if $n$ in (23) were a continuous variable.
This result is convenient because the factor $dk_{n}/dn$ will allow us to approximate sums over $n$ as integrals over $k$, at least in some ranges of $t$. As we will see, this mathematical convenience turns out to be what keeps the breathing of the trapped soliton looking just like the breathing of the untrapped soliton until times of the order of the trap period.
IV IV. Resuscitations
IV.1 IV.A The breathing mode
Figure 3: Condensate density in space and time, for an initially excited breathing mode in a harmonic trap with frequency $\omega$.
Figure 4 shows an example of the full density profile of the soliton breathing in a harmonic trap. Over many of its short periods, the mode is similar to that of the untrapped soliton, with a $1/\sqrt{t}$ dispersive decay in its amplitude. Over longer times, however—in particular after about one-half of the trap period—the amplitude increases again, showing periodic revivals, yet of an asymmetrical form. The behavior of the full density profile follows its peak at $x=0$ in a simple way, and so we focus on $|\psi(0,t)|$, which exhibits the trumpet-shaped asymmetrical resuscitations that we showed in Fig. 2.
For our small initial perturbation (5) from a perfect bright soliton, we have
$\psi(0,t)$ $=\Psi_{0}(0,t)+e^{i\frac{\hbar\kappa^{2}}{2M}t}\delta\psi(0,t)$
$=e^{i\frac{\hbar\kappa^{2}}{2M}t}\left(\frac{\kappa}{\sqrt{g}}+\delta\psi(0,t)\right)$
$\delta\psi(0,t)$ $=\sum_{n=0}^{\infty}\left(\beta_{n}e^{-i\Omega_{n}t}u_{n}(0)+\beta_{n}^{*}e^{i\Omega_{n}t}v_{n}^{*}(0)\right)$
$=\frac{2Z_{n}}{\sqrt{2\pi}}\sum_{n=0}^{\infty}\frac{\beta_{n}e^{-i\Omega_{n}t}k_{n}^{2}+\beta_{n}^{*}e^{i\Omega_{n}t}\kappa^{2}}{(k_{n}^{2}+\kappa^{2})}\;,$ (27)
when we use
$(u_{n},v_{n})=Z_{n}[(u_{k_{n}},v_{k_{n}})+(-1)^{n}(u_{-k_{n}},v_{-k_{n}})]$ (28)
and insert $u_{k}(0),v_{k}(0)$ from (II.2). The discrete $\beta_{n}$ are defined from the initial $\delta\psi(x,0)$ by projection onto the complete set of discrete $(u_{n},v_{n})$:
$\beta_{n}=\int\!dx\,\left(u^{*}_{n}(x)\delta\psi(x,0)-v^{*}_{n}(x)\delta\psi^{*}(x,0)\right)\;.$ (29)
With $\delta\psi(x,0)$ still given by our initial breathing perturbation (II.2), all $\beta_{n}$ with odd $n$ vanish by parity. Furthermore $\delta\psi(x,0)$ essentially vanishes outside the Soliton Zone, and so the integral over $x$ in (29) is the same as the analogous integral for $\beta_{k}$ in the trap-free case (II.3), just with $k\to k_{n}$ and an additional factor of $2Z_{n}$:
$\beta_{n}=-\sqrt{\frac{2\pi}{g}}\,\delta\,Z_{n}\mathrm{sech}\Big{(}\frac{\pi k_{n}}{2\kappa}\Big{)}$ (30)
for all even $n$, but $\beta_{n}=0$ for all odd $n$. Since $\beta_{n}$ are thus real, we obtain a compact sum over even $n$ for $|\psi(0,t)|^{2}$ to order $\delta$:
$|\psi(0,t)|^{2}$ $=\frac{\kappa^{2}}{g}[1-\delta\,B(t)+\mathcal{O}(\delta^{2})]$
$B(t)$ $=\frac{4}{\kappa}\sum_{n=0}^{\infty}Z_{2n}^{2}\mathrm{sech}\Big{(}\frac{\pi k_{2n}}{2\kappa}\Big{)}\cos(\Omega_{2n}t)\;.$ (31)
We identify this $B(t)$ as the trapped soliton’s breathing amplitude, analogous to the $B_{0}(t)$ for the untrapped soliton in (II.3).
To evaluate $B(t)$ explicitly, we combine our results (24,23,25,26) for $\Omega_{n}$, $k_{n}$, and $Z_{n}$, in the form
$B(t)$ $=2\mathrm{Re}\left(e^{i\frac{\hbar\kappa^{2}}{2M}t}e^{i\frac{\omega}{2}t}\sum_{n=0}^{\infty}\frac{1}{\kappa}\frac{dk_{2n}}{dn}\mathrm{sech}\Big{(}\frac{\pi k_{2n}}{2\kappa}\Big{)}e^{2i\omega t\left(n+\frac{1}{\pi}\cos^{-1}\frac{1-2\varepsilon^{2}n}{1+2\varepsilon^{2}n}\right)}\right)$
$=2\mathrm{Re}\left(e^{i\frac{\hbar\kappa^{2}}{2M}t}\sum_{n=0}^{\infty}\frac{1}{\kappa}\frac{dk_{2n}}{dn}\mathrm{sech}\Big{(}\frac{\pi k_{2n}}{2\kappa}\Big{)}e^{i\frac{\hbar k_{2n}^{2}}{2M}t}\right)\;.$ (32)
As long as each term in the sum varies slowly with $n$, the sum approaches a Riemann sum and we can use the Euler-Maclaurin series
$\sum_{n=0}^{\infty}f(\Delta n)=\frac{1}{\Delta}\int_{0}^{\infty}dxf(x)-\left.\frac{f(x)}{2}\right|_{0}^{\infty}+\mathcal{O}(\Delta^{2})$ (33)
to approximate (IV.1) as the $B_{0}(t)$ (II.3) that we found for the breathing amplitude of the untrapped soliton:
$B(t)=2\mathrm{Re}\left(e^{i\frac{\hbar\kappa^{2}}{2M}t}\int_{0}^{\infty}\!\frac{dk}{\kappa}\,\mathrm{sech}\Big{(}\frac{\pi k}{2\kappa}\Big{)}e^{i\frac{\hbar k^{2}}{2M}t}\right)+\mathcal{O}(\varepsilon)\;,$ (34)
where the factor of two appears instead of extending the lower limit of integration to -$\infty$, as in (II.3), since the integrand is even in $k$.
Do the terms in the sum (IV.1) actually vary slowly with $n$, though? Whenever $n$ appears in the summand (IV.1), it is indeed multiplied by the small factor $\varepsilon^{2}$—except for the one term $2i\omega tn$ that appears in the exponent. For times much less than the trap period, this term also varies slowly with $n$, and so for $t\ll 1/\omega$ the trapped bright soliton still breathes just as it does without any trap, and the breathing amplitude decays in a $1/t$ envelope. For times $t\gtrsim 1/\omega$, however, the sum (IV.1) is not well approximated by the integral (34). The non-Markovian nature of the Trap Zone as an environment for the breathing soliton becomes apparent.
For general times of order $t\sim 1/\omega$, the sum (IV.1) has no simple evaluation. We can consider times close to a whole number of half-periods, however,
$t=\frac{N\pi}{\omega}+\Delta t$ (35)
for whole-number $N$ and $|\Delta t|\ll 1/\omega$. For $N\gtrsim\varepsilon^{-2}$, $2N\cos^{-1}\big{(}(1-2\varepsilon^{2}n)/(1+2\varepsilon^{2}n)\big{)}$ does not vary slowly enough with $n$ for any integral approximation to be valid, but as long as $N\ll\varepsilon^{-2}$, the summand in (IV.1) again varies slowly with $n$, because $e^{2i\pi Nn}=1$. This yields the modified integral
$B(t)$ $=\mathcal{O}(\varepsilon)+2\mathrm{Re}\left(e^{i\frac{N\pi}{2}}e^{i\frac{\hbar\kappa^{2}}{2M}t}\int_{0}^{\infty}\!\frac{dk}{\kappa}\,\mathrm{sech}\Big{(}\frac{\pi k}{2\kappa}\Big{)}e^{i\frac{\hbar k^{2}}{2M}\Delta t}e^{2iN\cos^{-1}\big{(}\frac{\kappa^{2}-k^{2}}{k^{2}+\kappa^{2}}\big{)}}\right)=:B_{N}(\Delta t)+\mathcal{O}(\varepsilon)$ (36)
for
$B_{N}(\Delta t)$ $=2\mathrm{Re}\left[e^{i\frac{N\pi}{2}}e^{i\frac{\hbar\kappa^{2}}{2M}\Delta t}\int_{0}^{\infty}\!dz\,\mathrm{sech}\Big{(}\frac{\pi z}{2}\Big{)}e^{i\tau z^{2}}\left(\frac{1+iz}{1-iz}\right)^{2N}\right]\;,$ (37)
using $k\to\kappa z$ as the integration variable and noting that
$\cos^{-1}\left(\frac{1-z^{2}}{1+z^{2}}\right)\equiv i\ln\left(\frac{1-iz}{1+iz}\right)\;.$ (38)
After Eqns. (23-25) for the collective mode spectrum, Eqn. (37) for the resuscitation profiles in time is our paper’s final major result.
For $N=0$, $B_{0}(\Delta t)=B_{0}(-\Delta t)$, since $\Delta t\to-\Delta t$ just changes the term in square brackets in (37) to its complex conjugate and this does not affect its real part. For $N>0$, however, this no longer true, and $B_{N>0}(\Delta t)$ is asymmetric in $\Delta t$, with the asymmetry becoming more marked with increasing $N$. Figure 4 shows a direct comparison between $g|\psi(0,t)|^{2}/\tilde{\kappa}^{2}$ from Fig. 2 and the Bogoliubov-de Gennes approximations $(1+\delta)^{2}[1-\delta B_{N}(t-N\pi/\omega)]$ in intervals around three successive resuscitation times $t=N\pi/\omega$. Since $\varepsilon\doteq 0.14$ in this case is not actually very small, the integral approximation becomes poorer at larger $\Delta t$, and for later resuscitations it evidently exaggerates the severity of the post-revival drop in breathing amplitude. The agreement between nonlinear mean field theory and the Bogoliubov-de Gennes approximation is nevertheless close enough to confirm that the asymmetric profile of the resuscitations is indeed due to the slight non-uniformity of the collective mode frequency spectrum.
Figure 4: The same resuscitation sequence shown in Fig. 2 is here rendered faintly in the background, for the same $\delta=10^{-2}$ and $\varepsilon=(1+\delta)^{-2}/\sqrt{50}\doteq 0.14$. Superimposed in four time intervals around the revival times $t=N\pi/\omega$ are the Bogoliubov-de Gennes approximations $(1+\delta)^{2}[1-\delta B_{N}(t-N\pi/\omega)]$, where the integral approximation (37) has been used for $B_{N}(\Delta t)$, with the integrals evaluated numerically.
V V. Discussion
V.1 V.A Revivals
The initial attenuation of the soliton breathing mode within a $1/\sqrt{t}$ envelope is already expected from the case with no trapping potential. In the trapped case, it is then only natural to expect revivals after every trap half-period, once we realize that the attenuation of the excitation involves emission of particles, since a half-period is how long it takes before any initially emitted atoms will begin to return to the soliton. The revivals occur every half-trap-period, which is easy to deduce because our initial perturbation of the soliton is spatially even. In a trap, these even collective modes have eigenfrequencies $\Omega_{2n}$ separated by approximately $2\omega$, leading to a revival time estimate of $\pi/\omega$.
The less obvious phenomenon that we have seen in this paper, however, is that these expected breathing revivals—our “resuscitations”— have an asymmetric profile, which becomes more pronounced in later resuscitations. The breathing amplitude actually starts to revive distinctly earlier than the first $\pi/\omega$ half-period, and this early revival becomes steadily stronger in later resuscitations. Almost immediately after the half-period, then, the breathing amplitude falls sharply, before more gradually decaying until the next resuscitation.
V.2 V.B A qualitative explanation
Qualitatively, we can explain this phenomenon simply, from the spectral facts that we have described above. The two lowest modes of the Bogoliubov-de Gennes spectrum remain as the time-translation zero mode and the collective dipole mode with eigenfrequency exactly $\omega$, while all the other modes are raised above a large gap $\hbar\kappa^{2}/(2M)$, because excitations out of the condensate must all be excited above the large mean-field binding energy of the soliton. When the soliton is much narrower than the trap width, the spectrum above this chemical potential gap is still approximately like that of a single harmonic oscillator. Nevertheless the entire above-gap spectrum has to be slightly distorted away from the even level spacing of an oscillator, to “replace” the missing two lowest modes. These distortions are indeed supplied by the differences between the Bogoliubov-de Gennes equations and the Schrödinger equation of a single trapped particle.
In particular the slight systematic spectral distortion is such that the gaps between successive even-mode eigenfrequencies are consistently a bit greater than $2\omega$. This means that quasi-particles oscillate back and forth in the trap at slightly higher frequencies than $\omega$. Why are the atoms sped up in this way?
The atoms cannot oscillate faster than $\omega$ while they are in the Trap zone, outside the soliton’s mean field. While they are inside the soliton, however, the atoms evidently experience an effectively negative potential, giving them additional kinetic energy: higher speed. Although the soliton occupies only a narrow region in the middle of the trap, within this narrow region quasiparticles effectively experience quite a deep negative potential, and speed up significantly—as long as they are inside the soliton.
This effective burst of speed as quasiparticles pass through the soliton background lets them systematically return to the soliton again slightly earlier than they would in the trap with no soliton. Since the spectral stretching $d\nu/dn$ is not uniform, but affects lower eigenfrequencies more than higher ones in accordance with (25), there is a range of early return times. With every subsequent resuscitation, moreover, the lead of those faster quasiparticles steadily increases, in comparison to where they would be after the same amount of time without the speed boosts while they are inside the soliton. They return earlier and earlier every time, so that subsequent resuscitations start earlier and earlier.
The particles that return to the soliton before a full half-period keep moving on through the soliton, moreover, to continue oscillating back and forth in the trap. Shortly after the half-period, the early arrivals are already gone again, and the early amplification that they contributed to the revived breathing amplitude is now missing. This explains why the breathing amplitude drops after each half-period, and drops more and more noticeably in later resuscitations.
V.3 V.C The effect is not trivial
The asymmetric resuscitation profile should thus be comprehensible, given the spectral stretching that we have described in terms of speed boosts while within the soliton’s attraction. To improve upon this qualitative explanation, however, we have not been able to find any satisfactory analysis that is any simpler than the full Bogoliubov-de Gennes calculation that we have explained in our text and Appendices. The spectral distortion, and speeding-up that it implies for quasiparticles, is a non-trivial effect of coherent interaction between excited atoms and the background condensate.
The breathing excitations that we describe consist mostly of low-frequency quasiparticles (that is, with $\Omega_{n}$ close to the $\hbar\kappa^{2}/(2M)$ top of the mean-field gap). Outside the narrow background soliton, these quasiparticles are simply atoms, moving in a harmonic potential. Within the soliton background, however, these excited atoms are dressed by many-body effects: they become quasiparticles that involve many atoms. As quantum many-body effects go, these effects are still simple, inasmuch as they can be described within the linearized mean-field theory of Bogoliubov-de Gennes. Simpler than this, though, they are not.
The breathing mode of a bright soliton shakes particles out of the soliton. This means that a slightly imperfect bright soliton, with a breathing excitation, is an intrinsic atomic antenna, emitting atoms coherently with an inherent frequency that is defined, up to the limit of dispersive broadening, by the atomic interactions themselves. If the soliton breathes in a trap, the coherently emitted atoms return periodically. In what is perhaps one of the simplest possible examples of back-reaction, they interfere with their emitter itself. A bright soliton breathing in a trap is an intrinsic atomic interferometer.
Mean-field physics of dilute Bose-condensed gases is in general well understood, and so it is more complicated quantum many-body effects that may be of greater physical interest in experiments with trapped solitons. As a distinctive and non-trivial mean-field prediction, then, the asymmetrical resuscitations of the soliton breathing mode may be a useful background against which to look for subtler quantum effects, such as decay of the resuscitations due to decoherence from interactions among quasi-particles.
V.4 V.D Implications for non-Markovian evolution
Non-Markovian environments are in general so much more difficult to treat than Markovian ones that non-Markovian effects are often ignored, because even if they are not really small, they are often small in comparison to the effort required to compute them. If the open-system dynamics are represented in an equation of motion for system degrees of freedom alone, the memory effect of a non-Markovian environment typically makes this equation of motion integro-differential rather than differential. Integral equations of almost any kind are much harder to solve than purely differential equations.
Explicit results for non-Markovian open systems can sometimes be obtained by solving the evolution of the larger closed system that includes the environment—especially when this larger-system dynamics can be treated as linear, so that the non-Markovian evolution can be obtained by projection from a linear evolution of a large number of degrees of freedom. The present paper has been an example of this approach. And it has shown how even as simple a form of environment memory as oscillation in a weak harmonic trap can produce strange effects like our asymmetric resuscitations, when the simply non-Markovian environment interacts with a non-trivial open system such as the breathing mode of an atomic bright soliton.
Acknowledgement
This project traces its origins to T. Schlachter’s keen observation and early encouragement. We gratefully acknowledge his role in inspiring its initial direction.